The FraserRiver Company has purchased a new piece of factory equipment on January 1, 2018, and wishes to compare three depreciation methods: straight-line, double-declining-balance, and units-of-production The equipment costs $400,000 and has an estimated useful life of four years, or 8,000 hours. At the end of four years, the equipment is estimated to have a residual value of $20,000. Requirements 1. Use Excel to prepare depreciation schedules for straight-line, double-declining-balance, and units-of-production methods. Use cell references from the Data table. 2. Prepare a second depreciation schedule for double-declining-balance method, using the Excel function DDB. The DDB function cannot be used in the last year of the assets useful life. 3. At December 31, 2018, Fraser River is trying to determine if it should sell the factory equipment. Fraser River will only sell the factory equipment if the company earns a gain of at least $6,000. For each of the three depreciation methods, what is the minimum amount that Fraser River will sell the factory equipment for in order to have a gain of $6,000?
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

---Straight Line Method

$400,000 $20,000 4 2 Cost Residual Value Life (in vears 4 DepreciationAccumulated lexpense Ending Book Value Cost of Asset 18 $400,000 19$400,000 $400,000 $400,000 Depreciation 31-Dec- 31-Dec- 31-Dec-20 31-Dec-21 $95,000 $95,000 $95,000 $95,000 $95,000 $190,000 $285,000 $380,000 $305,000 $210,000 $115,000 $20,000 10

--Working with formula and cell reference

Cost of Asset

Depreciation expense

Accumulated Depreciation

Ending Book Value

400000

=($C$2-$C$3)/$C$4

=D7

=C7-E7

=+C7

=($C$2-$C$3)/$C$4

=E7+D8

=C8-E8

=+C8

=($C$2-$C$3)/$C$4

=E8+D9

=C9-E9

=+C9

=($C$2-$C$3)/$C$4

=E9+D10

=C10-E10

---Double Declining balance method

G34 12 13 14 15 16 17 18 19 Cost Residual Value Life (in vears Straight Line annual de Double declining rate $400,000 $20,000 4 $95,000 50% Beginning Double Declining Depreciation Accumulated 21 Book Value Rate expense DepreciationEnding Book Value 18 $400,000 31-Dec- 31-Dec-19 31-Dec-20 31-Dec-21 $200,000 $100,000 $50,000 $200,000 $100,000 $50,000 $30,000 $200,000 $300,000 $350,000 $380,000 $200,000 $100,000 $50,000 $20,000 23 24

--Working with formula and cell reference

Double declining rate

=(C17/(C14-C15))*2

Beginning Book Value

Double Declining Rate

Depreciation expense

Accumulated Depreciation

Ending Book Value

400000

=$C$18

=C22*D22

=E22

=C22-F22

=+G22

=$C$18

=C23*D23

=+F22+E23

=+C23-E23

=+G23

=$C$18

=C24*D24

=+F23+E24

=+C24-E24

=+G24

=50000-20000

=+F24+E25

=+C25-E25

---Units of production [assuming that 2000 hours are used each year]

B29 f Residual Value 27 28 29 30 31 Cost Residual Value Expected no. of hours $400,000 $20,000 8,000 Beginning Book Value lexpense Accumulated Depreciation Ending Book Value Depreciation 32 34 35 36 27 31-Dec- 31-Dec-19 31-Dec-20 31-Dec-21 18$400,000 $305,000 $210,000 $115,000 $95,000 $95,000 $95,000 $95,000 $95,000 $190,000 $285,000 $380,000 $305,000 $210,000 $115,000 $20,000

---Working

Beginning Book Value

Depreciation expense

Accumulated Depreciation

Ending Book Value

400000

=(($C$28-$C$29)/$C$30)*2000

=D33

=C33-E33

=+F33

=(($C$28-$C$29)/$C$30)*2000

=+E33+D34

=+C34-D34

=+F34

=(($C$28-$C$29)/$C$30)*2000

=+E34+D35

=+C35-D35

=+F35

=(($C$28-$C$29)/$C$30)*2000

=+E35+D36

=+C36-D36

  • Requirement 2

USING DDB Function

--Actual data and result

39 40 41 42 43 Cost Residual Value Life (in vears $400,000 $20,000 4 Accumulated Depreciation Beginning Ending Book Value Depreciation 45 Book Value exense 47 48 49 31-Dec-18 31-Dec-19 31-Dec-20 31-Dec-21 $400,000 $200,000 $100,000 $50,000 $200,000 $100,000 $50,000 $30,000 $200,000 $300,000 $350,000 $380,000 $200,000 $100,000 $50,000 $20,000

--Working with formula and cell reference

Beginning Book Value

Depreciation expense

Accumulated Depreciation

Ending Book Value

400000

=DDB($C$40,$C$41,$C$42,1)

=D46

=C46-E46

=+F46

=DDB($C$40,$C$41,$C$42,2)

=+E46+D47

=+C47-D47

=+F47

=DDB($C$40,$C$41,$C$42,3)

=+E47+D48

=+C48-D48

=+F48

30000

=+E48+D49

=+C49-D49

  • Requirement 3

Book Value at 31 Dec 2018

Gain required

Minimum amount at which the asset should be sold

Straight Line Method

$305,000

$6,000

$311,000

Double declining balance

$200,000

$6,000

$206,000

Units of production

$305,000

$6,000

$311,000